Instrument Log

10 MT IRT C1 Lamont, OK

Date

Operator

Comments

6/24/15

CAM

Data collections have failed for the past 5 hours, last successful at 1404 GMT. Ping test to the Moxa inside the IRT enclosure was successful. Connected laptop directly to data logger and was able to monitor the IRT with LoggerNet. Power cycled the Moxa at 1945 GMT. Also removed a wasp nest and several wasps from the enclosure. The following collection was successful at 2004 GMT. All missing data was recovered.

9/17/13

JEM

The 10M IRT needs to be replaced with a newer style sensor. Also DQPR 3834 indicates a bias between the 10M and 25M IRT. R/R the *10M IRT @1905 GMT. Updated the datalogger program and RTS @1930 GMT. *Old - 1029 (Model - KT 19.85) *New - 3350 (Model - KT 19.85 II)

2/26/13

CAW

The site lost power at approximately 0350 GMT on 2/26 due to a severe winter storm. Powered central cluster UPS off at 1730 GMT. Power was restored at 1920 GMT. Powered UPS on at 1945 GMT. Instrument RTS status O.K and data collections resumed once SDS operations were restored.

2/8/13

JEM

DQPR #3616 requests the10M IRT lens be replaced. Lowered the tower @1522 GMT and R/R the lens @1525 GMT. Raised the tower @1527 GMT and RTS @1537 GMT.

12/19/12

JEM

DQPR #3616 requests that the10M tower be lowered and the 10M IRT lens be cleaned. Lowered the 10M tower @1625 GMT. Cleaned the lens @1630 GMT. RTS @1635 GMT.

10/31/12

JEM

The 10M IRT serial number needs to be verified. Lowered the tower and verified the serial number as 1029 GMT.

2/7/12

JEM

The tower needs to be lowered to measure the 10M MFR head cable length. Dropped the tower @2055 GMT. Measured the cable length and put the tower back up @2110 GMT.

2/6/12

JEM

The 10M MFR was upgraded from 1700 GMT to 2330 GMT. IRT power may have been disconnected at various times during the upgrade period. The tower was lowered from 2145 GMT to 2315 GMT. 10M IRT data will be affected during this time.

1/4/12

CAM

DQPR 2811 states that the IRT temperature has dropped and is very noisy. The logger temp and battery voltage are dropping out as well. Monitored battery voltage locally at data logger and found it to be steady at 13.7V. Tightened data logger power connections and power cycled logger at 1510 GMT. Noticed the data logger had no earth ground. Returned later and reseated the data logger to wiring panel connection at 2030 GMT and grounded the logger to earth ground at 2110 GMT. Data appeared to return to normal as viewed later on DQ Health and Status page.

4/15/11

JEM

The instrument has not collected since the 0400 hour. Due to other instruments that are down the source of the problems seems to be the Rocketport hub. The Rocketport power light is flashing. Power cycling the Rocketport hub did not correct the issue. Notified the IT department of the issue.

8/17/10

8/18/10

CAW

10M MFR is due for calibration. Lowered tower and removed MFR head at 1932 GMT on 8/17. Raised tower at 1935 GMT. MFR was calibrated and then reinstalled on 8/18 at 1755 GMT. RTS at 1759 GMT.

4/11/06

MK

The 10M IRT is not communicating with site data system. R/R bad optical bit driver located in the SDS facility.

12/5/05

MK

The OS needs to be upgraded on all CR10X's. Downloaded current data off of data logger, retrieved program from data logger, loaded new OS. Loaded program back into CR10X datalogger. System is operating normal. S/N 26908. RTS at 1915 GMT.

9/30/04

JEM

Collected data from the datalogger @1905 GMT due to a collection problem with the SDS. RTS @1923 GMT.

12/22/03

MR

The instrument mentor wants the lens checked and cleaned. Lowered the tower at 2025 GMT. Cleaned the lens. Raised the tower at 2033 GMT. RTS at 2040 GMT.

8/19/03

MR

The tower needs to be lowered due to maintenance on other equipment. Lowered the tower at 1615 GMT. Raised the tower at 1730.

2/4/03

JEM

The entire Central Facility has experienced a power outage at approximately 1845 GMT. Power was restored to the Central Facility at approximately 1945 GMT. RTS @2000 GMT.

10/4/01

MR

The tower needs to be lowered due to maintenance on other equipment. Lowered the tower at 1800 GMT. Raised the tower at 1820 GMT. RTS at 1830.

10/4/01

MR

The 10M tower needs to be lowered due to calibration of the 10M MFR. Lowered the tower at 1420 GMT. Raised the tower at 1500 GMT.

10/3/01

MR

The 10 meter tower needs to be lowered for calibration on other equipment. Lowered the tower at 2210 GMT. Raised the tower at 2225 GMT.

9/26/01 9/27/01

V. Morris

Moved IRT/10m signal from MFR data logger to CSI data logger at 2100 GMT on 9/26. Returned to site on 9/27 and reconfigured IRT and started CSI data logger at 1445 GMT. Cleaned IRT lens at 1450 GMT.

2/7/01

MR

Tower lowered at 1600 GMT for MFR maintenance. Raised the tower at 1605 GMT.

12/28/99

MK

Due to power line maintenance starting at 1500 GMT all instruments will be down unless battery power is present. At 1536 GMT the AC power was restored to all CF instruments. RTS at 1536 GMT.

9/2/99

ZW

Tower lowered at 1942 GMT for MFR maintenance. Site Ops raised the tower and RTS at 2003 GMT.

5/4/99

JT

The 10 Meter tower was replaced and the 10M IRT returned to service.

4/14/99

CW

The 10M tower was taken down at 2120 GMT for tower replacement/upgrade. 10M IRT was taken off line and will remain off line until tower replacement is complete.

12/24/98

CM/GS

Victor Morris requested that Site Ops fabricate some type of insulation to insulate the 10M IRT during the cold winter months. In the past the IRT data has been degraded when temperatures drop below freezing. Lowered the 10M tower at 1525 GMT. Removed the IRT and IRT mounting assembly to the lab to fabricate an insulation cover. Used 5/8" insulation material and aluminum tape to cover the 4 exposed sides of the IRT. Reformed the original aluminum housing to cover the insulation. Remounted the IRT, raised the tower and RTS at 1845 GMT.

 8/6/98

 GRS

Lowered the 10m tower at 1842 GMT to read some program values on the 10m IRT. The values in the OUTA mode was: T.low= -20.0 T.end= 180.0 Tr= 3.0 VOLT= 10. Raised the tower and RTS the instrument at 1908 GMT.